# VRBench: A Benchmark for Multi-Step Reasoning in Long Narrative Videos

### v1

@@ Abstract
We present VRBench, the first long narrative video benchmark crafted for evaluating large models’ multi-step reasoning capabilities, addressing limitations in existing evaluations that overlook temporal reasoning and procedural validity. It comprises 1,010 long videos (with an average duration of 1.6 hours), along with 9,468 human-labeled multi-step question-answering pairs and 30,292 reasoning steps with timestamps. These videos are curated via a multi-stage filtering process including expert inter-rater reviewing to prioritize plot coherence. We develop a human-AI collaborative framework that generates coherent reasoning chains, each requiring multiple temporally grounded steps, spanning seven types (e.g., event attribution, implicit inference). VRBench designs a multi-phase evaluation pipeline that assesses models at both the outcome and process levels. Apart from the MCQs for the final results, we propose a progress-level LLM-guided scoring metric to evaluate the quality of the reasoning chain from multiple dimensions comprehensively. Through extensive evaluations of 12 LLMs and 16 VLMs on VRBench, we undertake a thorough analysis and provide valuable insights that advance the field of multi-step reasoning.
Figure 1: Overview of VRBench. We present VRBench, a long narrative video benchmark for multi-step reasoning. VRBench includes 1,010 manual-filtered narrative videos, covering 8 languages and 7 video categories that are suitable for reasoning about temporal relations. We also provide high-quality stepwise annotations for reasoning, which are labeled and reviewed by human experts. Each video incorporates 8-10 complex question-answer pairs, multi-step reasoning chain, and fine-grained timestamps. To fully evaluate the capability of models in multi-step reasoning, we propose a multi-phase evaluation pipeline that assesses model results both from the process- and outcome-level. Our VRBench is the first video reasoning benchmark that both supports multi-step annotation and evaluation.
@@ 1 Introduction
In data construction, VRBench aggregates 1,010 meticulously selected videos (1.6h average duration) across seven narrative categories (e.g., movies, sports, travelogues), sourced through a multi-stage filtering process. Our curation pipeline combines automated retrieval with expert validation (inter-rater reliability $\rho$=0.82), prioritizing plot coherence over domain-specific knowledge. This contrasts with existing benchmarks like MMVU [89] that emphasize disciplinary competence. To conduct multi-step reasoning chain annotation, we develop a human-AI collaborative approach, generating 8-10 QA pairs per video with explicit temporal grounding. Each question requires no less than 2 reasoning steps annotated with precise timestamps (Figure 3), validated through iterative expert review (95% inter-annotator agreement). The taxonomy spans seven reasoning types from event prediction to implicit inference, ensuring comprehensive coverage of narrative analysis skills.
When evaluating models on VRBench, we propose a multi-phase evaluation pipeline that combines outcome verification with process analysis. Beyond conventional multiple-choice accuracy that evaluates the outcome-level performance, we further introduce the process-level metric LLM-guided scoring to evaluate the overall reasoning process quality. This dual approach reveals critical insights – for instance, while GPT-4o achieves 83.25% outcome accuracy, its process rating scores drop to 58.1%, exposing reasoning fragility. Evaluations of 28 state-of-the-art large models reveal proprietary VLMs with long-context support outperform text-only LLMs by 12.2% absolute, emphasizing the importance of dense visual grounding. System-2 optimization strategies yield disproportionate gains in process metrics, and despite parameter parity, open-source VLMs lag behind proprietary counterparts by 8.99%, suggesting architectural limitations beyond scale. Our ablation studies show strong alignment between human and LLM evaluations and quantify the impact of test-time scaling—doubling context windows improves QwQ-32B’s [62] accuracy by 12.43%.
@@ 3.1 Video Curation
Upon the aforementioned two quantitative prerequisites, we organize a panel of 14 multilingual domain experts to evaluate candidate videos using a standardized 10-point scale based on plot coherence and richness. The scoring is guided by several instructions. Videos scoring below 7 are systematically excluded, resulting in a final curated set of 1,010 high-quality narratives. Detailed annotation protocols are documented in the supplementary materials.
@@ 3.2 Stepwise Reasoning Annotations
We recruit and train 67 graduate students to generate 8-10 high-quality QA pairs per video, providing raw video footage, translated subtitles, and GPT-generated pre-annotations that offer contextual hints without meeting final benchmark standards. Each QA pair must satisfy four rigorous criteria: (1) Non-synopsis questions require $\geq$2 timestamped reasoning steps (start/end times documented); (2) Temporal distribution constraints ($\leq$4 questions from 0-15min, $\geq$3 from 15-40min, $\geq 1$ from 40-120min); (3) Coverage of $\geq$5 predefined reasoning taxonomies (from 7 categories) ensuring diversity; (4) Mandatory multimodal grounding where solutions demand both visual analysis (excluding subtitle-only answers) and explicit reasoning (beyond basic perception).
@@ 3.3 Multi-Phase Evaluation Pipeline
We benchmark VLMs and LLMs through a comprehensive multi-phase evaluation pipeline, which compares predictions against ground-truth annotations both at the process-level and outcome level. For the outcome-level evaluation stage, we adopt a multiple-choice question (MCQ) format, where false options are generated by DeepSeek-V3 [38] using carefully-designed prompts based on human-annotated answers. For the process-level stage, we propose the open-ended rating to fully evaluate the model’s multi-step reasoning capability. Specifically, we adopt an LLM to evaluate the overall quality of the whole reasoning process through four 0-10 scores: logical coherence (40%), similarity to ground truth (40%, excluded for event prediction and hypothetical reasoning tasks), factual accuracy (10%), and clarity (10%). DeepSeek-V3 [38] serves as the judge due to its optimal balance between cost and human alignment (Section 4). For the event prediction and hypothetical reasoning tasks, we argue that there is no ground-truth reasoning steps since there might be multiple possible predictions, hence, we remove the similarity score and compute the final score with an 8:1:1 weight. All metrics are normalized to 0-100 scales, and we report the rankings by computing average scores across all metrics.
@@
[…]
Table 2: Evaluation results of LLMs and VLMs on VRBench across three evaluation metrics (outcome-level MCQ, process-level MCQ, process-level open-ended evaluation) and seven QA taxonomies (event attribution, counting problems, hypothetical reasoning, implicit inferences, information synopsis, event prediction, logical linkage). bold values indicate the best results among all models, and Underlined values are best results of LLMs.
@@ 4.1 Main Results
Table 2 shows the CoT evaluation results of LLMs and VLMs on VRBench. Among LLMs, we found the proprietary model Gemini-2.0-Flash-Thinking [12] achieves the optimal performance while other reasoning models such as OpenAI o1-preview [48] and DeepSeek-R1 [18] also demonstrate strong results. It is noted that the open-source model DeepSeek-R1 [18] achieves the best outcome-level MCQ accuracy of 70.26%. For the VLMs, Gemini-2.0-Pro [60] achieves 76.61% overall accuracy, surpassing all components by at least 5.93%. GPT-4o [47] and Claude-3.7-Sonnet [59] also show competitive results, which demonstrate 70.68% and 70.17% overall performance, respectively. For open-source models, InternVL2.5-78B [7], emerges as the best non-proprietary VLMs with a 66.10% overall accuracy.
@@ 4.2 Ablations and Analysis
The role of System-2 thinking. o1-like LLMs, also denoted as system-2 models, have emerged with superior reasoning abilities in multi-disciplinary scenarios such as science and math. We further explore these models’ performance on the narrative video benchmark. We both assess the role of open-source and proprietary o1-models, as well as the gap compared with their previous system-1 version. For proprietaries, we compare OpenAI o1-preview [48] with GPT-4o [47], and results show that the system-2 version o1 outperforms GPT-4o with 4.30% overall accuracy. For open-source model DeepSeek-V3 [38] and R1 [11], we observe that DeepSeek-R1 achieves higher scores both on multiple-choice questions and open-ended evaluations. It is noted that compared with Qwen2.5-72B-Instruct [79], QwQ-32B [62] gets higher reasoning ratings yet lower outcome-level MCQ results. This shows that though some system-2 models are more proficient in multi-step thinking, their ability to converge lengthy thinking processes into correct answers still needs improvement. Recently, some o1-like training strategies have also been utilized by some VLMs, such as LongVA-7B-DPO [88], a model that utilizes direct preference optimization [52] for long video understanding. It achieves a 61.44% overall performance and surpasses its vanilla version, suggesting that training VLMs with different optimization methods is feasible and developing system-2 VLM models is crucial for tackling multi-step reasoning questions in narrative videos.
@@
Impact of Model Size. It is commonly believed that models with large scales are often accompanied by better perception and reasoning capabilities. To verify such claims, we conduct experiments on the models with different scales, such as Qwen2.5-7B [80] and Qwen2.5-72B, Qwen2.5-VL [2] with 7B and 72B parameters, and InternVL2.5-8B [7] and InternVL2.5-78B. As shown in Table 2, models with more parameters obtain higher overall accuracy (50.70% vs 56.18% for Qwen2.5, 63.06% vs 69.23% for Qwen2.5-VL, and 56.11% vs 66.10% for InternVL2.5), and the performance gap on process-level metrics is more apparent. This indicates that large-scale models are more likely to achieve advanced reasoning capabilities. However, we also noticed that models specifically trained on the reasoning corpus like QwQ-32B [62], can achieve comparable performance compared with its 72B Qwen2.5 counterpart (55.96% vs. 56.18% in overall accuracy) with a smaller parameter amount. This suggests that training small-size models with various preference optimization strategies and reasoning-centric corpora might improve its overall reasoning capability.
@@
Long Context Helps. We observe a substantial performance gap between models with and without long context support. As shown in Table 2, models with long frame input tend to achieve higher overall accuracy. For example, Gemini-2.0-Pro [60] is capable of inference with 0.5 fps and supports large number frame inputs when answering questions of videos over 1 hour, and it achieves the optimal performance compared with all other fixed-length models. For the open-source models, Qwen2.5-VL-7B [2] and VideoChat-Flash [32] with 512 frame input achieve 63.06% and 57.30% overall accuracy, which is 6.95% and 1.19% higher than the 8B InternVL-2.5 [7] model with 64 frames input. Since the average length of VRBench is 1.61 hours, the necessity of long context support is becoming more pronounced, so that models can perceive more plot elements that help the analysis of long video content.
@@ 4.3 Test-Time Scaling Exploration
We here probe models’ performance when scaling test-time compute cost. Since we tend to prompt the models with a Chain-of-Thought template when tackling complex reasoning tasks, it is observed that system-2 models are capable of dramatically expanding the inference budget to improve their performance. To this end, we set a series of token limitations on these models and require the models to think with different instructions. Specifically, we select QwQ-32B [62] as the experimental LLM and Qwen2-VL-7B [79] as the testing VLM on a subset with 300 videos and 2,403 questions. For each model, several parallel experiments are conducted with increasing maximum token limits from 256 to 2048. We instruct the models to output the reasoning process as much as possible for large token limitations, and let them think concisely when the token limit is low. Since we aim to investigate the quality of models’ initial reasoning process and their capability to generate the right answers through thinking, we report the overall scores that both evaluate the outcome accuracy and process quailty. Results are shown in Figure 5, where QwQ shows a remarkable rating boost from 48.91% to 61.34% when setting a large token limit. On the contrary, the small-sized system-1 model Qwen2-VL-7B performs even worse when using a large token number limitation and instructing the model to output more thinking process. The model tends to output more ambiguous outputs that lead to the wrong answer. This observation suggests that models with large parameter scales and system-2 capabilities benefit from the test-time scaling strategy. It also provides the insight that developing more test-time scaling approaches that guide the model to generate long reasoning traces could be a feasible and promising way to unlock more potential capabilities of system-2 models, thereby benefiting complex tasks that require multi-step reasoning.

### v2

@@ Abstract
We present VRBench, the first long narrative video benchmark crafted for evaluating large models’ multi-step reasoning capabilities, addressing limitations in existing evaluations that overlook temporal reasoning and procedural validity. It comprises 960 long videos (with an average duration of 1.6 hours), along with 8,243 human-labeled multi-step question-answering pairs and 25,106 reasoning steps with timestamps. These videos are curated via a multi-stage filtering process including expert inter-rater reviewing to prioritize plot coherence. We develop a human-AI collaborative framework that generates coherent reasoning chains, each requiring multiple temporally grounded steps, spanning seven types (e.g., event attribution, implicit inference). VRBench designs a multi-phase evaluation pipeline that assesses models at both the outcome and process levels. Apart from the MCQs for the final results, we propose a progress-level LLM-guided scoring metric to evaluate the quality of the reasoning chain from multiple dimensions comprehensively. Through extensive evaluations of 12 LLMs and 19 VLMs on VRBench, we undertake a thorough analysis and provide valuable insights that advance the field of multi-step reasoning.
Figure 1: Overview of VRBench. We present VRBench, a long narrative video benchmark for multi-step reasoning. VRBench includes 960 manual-filtered narrative videos, covering 8 languages and 7 video categories that are suitable for reasoning about temporal relations. We also provide high-quality stepwise annotations for reasoning, which are labeled and reviewed by human experts. Each video incorporates 8-10 complex question-answer pairs, a multi-step reasoning chain, and fine-grained timestamps. To fully evaluate the capability of models in multi-step reasoning, we propose a multi-phase evaluation pipeline that assesses model results both from the process and outcome levels. Our VRBench is the first video reasoning benchmark that supports both multi-step annotation and evaluation.
@@ 1 Introduction
In data construction, VRBench aggregates 960 meticulously selected videos (1.6h average duration) across seven narrative categories (e.g., movies, sports, travelogues), sourced through a multi-stage filtering process. Our curation pipeline combines automated retrieval with expert validation (inter-rater reliability $\rho$=0.82), prioritizing plot coherence over domain-specific knowledge. This contrasts with existing benchmarks like MMVU [92] that emphasize disciplinary competence. To conduct multi-step reasoning chain annotation, we develop a human-AI collaborative approach, generating 8-10 QA pairs per video with explicit temporal grounding. Each question requires no less than 2 reasoning steps annotated with precise timestamps (Figure 3), validated through iterative expert review (95% inter-annotator agreement). The taxonomy spans seven reasoning types from event prediction to implicit inference, ensuring comprehensive coverage of narrative analysis skills.
When evaluating models on VRBench, we propose a multi-phase evaluation pipeline that combines outcome verification with process analysis. Beyond conventional multiple-choice accuracy that evaluates the outcome-level performance, we further introduce the process-level metric LLM-guided scoring to evaluate the overall reasoning process quality. This dual approach reveals critical insights – for instance, while GPT-4o achieves 81.25% outcome accuracy, its process rating scores drop to 56.11%, exposing reasoning fragility. Evaluations of 31 state-of-the-art large models reveal proprietary VLMs with long-context support outperform text-only LLMs by 13.82% absolute, emphasizing the importance of dense visual grounding. System-2 optimization strategies yield disproportionate gains in process metrics, and despite parameter parity, open-source VLMs lag behind proprietary counterparts by 12.30%, suggesting architectural limitations beyond scale. Our ablation studies show strong alignment between human and LLM evaluations and quantify the impact of test-time scaling—doubling context windows improves QwQ-32B’s [65] accuracy by 12.43%.
@@ 3.1 Video Curation
Upon the aforementioned two quantitative prerequisites, we organized a panel of 14 multilingual domain experts to evaluate candidate videos using a standardized 10-point scale based on plot coherence and richness. The scoring is guided by several instructions. Videos scoring below 7 are systematically excluded, resulting in a final curated set of 960 high-quality narratives. Detailed annotation protocols are documented in the supplementary materials.
@@ 3.2 Stepwise Reasoning Annotations
We recruit and train 67 graduate students to generate 8-10 high-quality QA pairs per video, providing raw video footage, translated subtitles, and GPT-generated pre-annotations that offer contextual hints without meeting final benchmark standards. Each QA pair must satisfy four rigorous criteria: (1) Non-synopsis questions require $\geq$2 timestamped reasoning steps (start/end times documented); (2) Temporal distribution constraints ($\leq$4 questions from 0-15min, $\geq$3 from 15-40min, 1$\geq$ 1≥ 1 from 40-120min); (3) Coverage of $\geq$5 predefined reasoning taxonomies (from 7 categories) ensuring diversity; (4) Mandatory multimodal grounding where solutions demand both visual analysis (excluding subtitle-only answers) and explicit reasoning (beyond basic perception).
@@ 3.3 Multi-Phase Evaluation Pipeline
We benchmark VLMs and LLMs through a comprehensive multi-phase evaluation pipeline, which compares predictions against ground-truth annotations both at the process and outcome level. For the outcome-level evaluation stage, we adopt a multiple-choice question (MCQ) format, where false options are generated by DeepSeek-V3 [38] using carefully-designed prompts based on human-annotated answers. For the process-level stage, we propose the open-ended rating to fully evaluate the model’s multi-step reasoning capability. Specifically, we adopt an LLM to evaluate the overall quality of the whole reasoning process through four 0-10 scores: logical coherence (40%), similarity to ground truth (40%, excluded for event prediction and hypothetical reasoning tasks), factual accuracy (10%), and clarity (10%). DeepSeek-V3 [38] serves as the judge due to its optimal balance between cost and human alignment (Section 4). For the event prediction and hypothetical reasoning tasks, we argue that there is no ground-truth reasoning steps since there might be multiple possible predictions, hence, we remove the similarity score and compute the final score with an 8:1:1 weight. All metrics are normalized to 0-100 scales, and we report the rankings by computing average scores across all metrics.
@@
Model Overall Results by Metric Results by Taxonomy MCQ-O OE-P EA CP HR II IS EP LL LLMs Proprietary Models GPT-4o [47] 55.49 63.87 47.11 54.41 32.63 69.03 61.47 71.83 66.37 66.19 Claude-3.7-Sonnet [59] 58.07 62.91 53.23 55.87 35.11 69.83 63.29 72.61 67.91 67.89 o1-preview [48] 60.14 68.47 51.81 56.97 35.87 70.51 64.81 73.13 68.27 69.41 Gemini-2.0-Flash-Thinking [61] 60.97 67.53 54.41 58.61 38.11 72.09 67.57 74.19 70.43 71.13 Open-Source Models QwQ-32B-preview [64] 35.90 27.51 44.29 34.31 34.41 46.29 39.67 27.97 42.39 40.61 InternLM3-8B-Instruct [3] 47.81 50.31 45.31 45.47 34.87 60.33 51.77 55.47 56.97 56.81 Qwen2.5-7B-Instruct [83] 48.29 52.61 43.97 46.93 35.17 61.83 51.57 54.87 56.63 57.03 Llama3.3-70B-Instruct [41] 49.84 52.59 47.09 47.63 38.07 63.87 54.03 54.83 59.47 56.97 QwQ-32B [65] 52.52 56.01 49.03 49.53 36.03 62.01 54.03 52.03 58.01 59.03 Qwen2.5-72B-Instruct [83] 53.51 60.49 46.53 51.03 36.01 67.03 58.53 67.83 61.93 63.03 DeepSeek-V3 [38] 56.06 64.79 47.33 54.03 36.57 69.97 61.83 69.53 65.53 65.43 DeepSeek-R1 [18] 57.13 64.19 50.07 56.13 38.91 72.81 64.13 68.01 68.53 67.93 VLMs Proprietary Models Claude-3.7-Sonnet [59] 68.15 80.09 56.21 63.11 32.97 72.63 71.13 75.37 71.33 70.27 GPT-4o [47] 68.68 81.23 56.13 66.61 36.51 76.67 70.47 78.03 72.03 73.17 Gemini-2.0-Pro [61] 74.61 83.29 65.93 71.09 65.21 81.01 75.73 87.13 77.93 75.89 Open-Source Models DeepSeek-VL2 [79] 31.50 33.27 29.73 25.93 22.41 35.73 31.73 30.01 33.29 29.57 H2OVL Mississippi-2B [17] 47.15 52.33 41.97 35.17 32.37 51.71 41.41 60.03 49.97 41.91 Phi-3.5-Vision [1] 48.52 58.03 39.01 31.53 28.03 45.03 37.03 68.03 44.03 36.03 LongVA-7B [91] 50.14 67.81 32.47 27.61 25.07 38.69 33.77 76.91 38.27 32.07 InternVL2.5-8B [7] 50.41 69.31 31.51 26.63 26.31 37.99 31.21 85.97 37.37 30.69 MiMo-VL-7B-RL [60] 50.48 63.39 37.57 47.31 34.59 63.83 53.23 74.61 60.33 56.97 VideoChat-Flash-7B [32] 50.82 72.01 29.63 24.69 21.41 39.37 29.17 79.37 34.87 28.07 InternVideo2.5 [73] 51.94 75.63 28.25 23.81 24.09 34.51 27.07 84.57 33.61 26.81 LongVA-7B-DPO [91] 52.36 67.91 36.81 30.73 27.17 45.09 38.23 79.39 43.67 37.03 Qwen2-VL-7B [68] 54.08 72.01 36.15 30.49 26.61 46.71 33.69 85.27 43.31 35.29 Aria [30] 54.55 72.97 36.13 30.19 29.23 44.23 34.63 86.23 44.99 34.57 Qwen2.5-VL-7B [2] 56.52 69.61 43.43 37.07 33.17 54.27 42.27 83.91 50.17 42.87 Keye-VL-8B-Preview [63] 60.44 64.41 56.47 62.53 40.37 73.61 67.99 70.13 71.83 71.03 Qwen2.5-VL-72B [2] 61.71 66.85 56.57 51.87 46.13 67.13 54.13 90.04 63.67 60.77 Kimi-VL-A3B-Thinking-2506 [62] 61.82 61.67 61.97 64.53 47.91 71.37 69.57 74.47 71.23 71.47 InternVL2.5-78B [7] 62.31 76.61 48.01 43.77 38.67 58.21 46.13 87.53 54.43 47.57
Table 2: Evaluation results on VRBench across two evaluation metrics (outcome-level MCQ, process-level open-ended evaluation) and seven QA taxonomies (event attribution, counting problems, hypothetical reasoning, implicit inferences, information synopsis, event prediction, logical linkage). Bold values indicate the best results among all models, and underlined values are best results of LLMs.
@@ 4.1 Main Results
Table 2 shows the CoT evaluation results of LLMs and VLMs on VRBench. Among LLMs, we found the proprietary model Gemini-2.0-Flash-Thinking [12] achieves the optimal performance while other reasoning models such as OpenAI o1-preview [48] and DeepSeek-R1 [18] also demonstrate strong results. It is noted that o1-preview [48] achieves the best outcome-level MCQ accuracy of 68.47%. For the VLMs, Gemini-2.0-Pro [61] achieves 74.61% overall accuracy, surpassing all components by at least 5.93%. GPT-4o [47] and Claude-3.7-Sonnet [59] also show competitive results, which demonstrate 68.68% and 68.15% overall performance, respectively. For open-source models, InternVL2.5-78B [7], emerges as the best non-proprietary VLMs with a 62.31% overall accuracy.
@@ 4.2 Ablations and Analysis
The role of System-2 thinking. o1-like LLMs, also denoted as system-2 models, have emerged with superior reasoning abilities in multi-disciplinary scenarios such as science and math. We further explore these models’ performance on the narrative video benchmark. We both assess the role of open-source and proprietary o1-models, as well as the gap compared with their previous system-1 version. For proprietaries, we compare OpenAI o1-preview [48] with GPT-4o [47], and results show that the system-2 version o1 outperforms GPT-4o with 4.30% overall accuracy. For open-source model DeepSeek-V3 [38] and R1 [11], we observe that DeepSeek-R1 achieves higher scores both on multiple-choice questions and open-ended evaluations. It is noted that compared with Qwen2.5-72B-Instruct [82], QwQ-32B [65] gets higher reasoning ratings yet lower outcome-level MCQ results. This shows that though some system-2 models are more proficient in multi-step thinking, their ability to converge lengthy thinking processes into correct answers still needs improvement. Recently, some o1-like training strategies have also been utilized by some VLMs, such as LongVA-7B-DPO [91], a model that utilizes direct preference optimization [52] for long video understanding. It achieves an overall performance of 52.36% and surpasses its vanilla version, suggesting that training VLMs with different optimization methods is feasible and that developing system-2 VLM models is crucial for tackling multi-step reasoning questions in narrative videos.
@@
Impact of Model Size. It is commonly believed that models with large scales are often accompanied by better perception and reasoning capabilities. To verify such claims, we conduct experiments on the models with different scales, such as Qwen2.5-7B [83] and Qwen2.5-72B, Qwen2.5-VL [2] with 7B and 72B parameters, and InternVL2.5-8B [7] and InternVL2.5-78B. As shown in Table 2, models with more parameters obtain higher overall accuracy (48.29% vs 53.51% for Qwen2.5, 56.52% vs 61.71% for Qwen2.5-VL, and 50.41% vs 62.31% for InternVL2.5), and the performance gap on process-level metrics is more apparent. This indicates that large-scale models are more likely to achieve advanced reasoning capabilities. However, we also noticed that models specifically trained on the reasoning corpus like QwQ-32B [65], can achieve comparable performance compared with its 72B Qwen2.5 counterpart (52.52% vs. 53.51% in overall accuracy) with a smaller parameter amount. This suggests that training small-sized models with various preference optimization strategies and reasoning-centric corpora might improve their overall reasoning capability.
@@
Long Context Helps. We observe a substantial performance gap between models with and without long context support. As shown in Table 2, models with long frame input tend to achieve higher overall accuracy. For example, Gemini-2.0-Pro [61] is capable of inference with 0.5 fps and supports large number frame inputs when answering questions of videos over 1 hour, and it achieves the optimal performance compared with all other fixed-length models. For the open-source models, Qwen2.5-VL-7B [2] and Kimi-VL-A3B-Thinking-2506 [62] with 256 frame input achieve 61.71% and 61.82% overall accuracy, which is 11.30% and 11.41% higher than the 8B InternVL-2.5 [7] model with 64 frames input. Since the average length of VRBench is 1.61 hours, the necessity of long context support is becoming more pronounced, so that models can perceive more plot elements that help the analysis of long video content.
@@ 4.3 Test-Time Scaling Exploration
We here probe models’ performance when scaling test-time compute cost. Since we tend to prompt the models with a Chain-of-Thought template when tackling complex reasoning tasks, it is observed that system-2 models are capable of dramatically expanding the inference budget to improve their performance. To this end, we set a series of token limitations on these models and require the models to think with different instructions. Specifically, we select QwQ-32B [65] as the experimental LLM and Qwen2-VL-7B [82] as the testing VLM on a subset with 300 videos and 2,403 questions. For each model, several parallel experiments are conducted with increasing maximum token limits from 256 to 2048. We instruct the models to output the reasoning process as much as possible for large token limitations, and let them think concisely when the token limit is low. Since we aim to investigate the quality of models’ initial reasoning process and their capability to generate the right answers through thinking, we report the overall scores that both evaluate the outcome accuracy and process quality. Results are shown in Figure 5, where QwQ shows a remarkable rating boost from 48.91% to 61.34% when setting a large token limit. On the contrary, the small-sized system-1 model Qwen2-VL-7B performs even worse when using a large token number limitation and instructing the model to output more thinking process. The model tends to output more ambiguous outputs that lead to the wrong answer. This observation suggests that models with large parameter scales and system-2 capabilities benefit from the test-time scaling strategy. It also provides the insight that developing more test-time scaling approaches that guide the model to generate long reasoning traces could be a feasible and promising way to unlock more potential capabilities of system-2 models, thereby benefiting complex tasks that require multi-step reasoning.